ganeshie8 (ganeshie8):

Show that if Ax is in the nullspace of A^T, then Ax=0

jhonyy9 (jhonyy9):

@ganeshie8 details about Ax and T ? please ---- ty.

OpenStudy (empty):

For a vector \(v\) to be in a matrix \(B\)'s null space it must satisfy this equation by definition: \[Bv=0\] Similarly, \(Ax\) is a vector, since a matrix times a vector is a vector, and for it to be in the nullspace of \(A^\top\) that means: \[A^\top (Ax) = 0\] So far just definitions. The trick that took me a few minutes to figure out it, what's the length of the vector \(Ax\)?

OpenStudy (p0sitr0n):

\[Ker(A^T)=Im(A)^{orthogonal}\]

ganeshie8 (ganeshie8):

Awesome! So that states kernel of A^T is orthogonal to column space So if a vector exists in both column space and the kernel of A^T, then it must be the zero vector. Ker(A^T) \(\cap\) Im(A) = zero vector